For the following reaction, 14.8 grams of sulfur are allowed to react with with 19.9 grams of carbon monoxide.

sulfur (s) + carbon monoxide (g) sulfur dioxide (g) + carbon (s)

What is the maximum amount of sulfur dioxide that can be formed?
1
Expert's answer
2013-02-15T04:37:13-0500
S + 2CO = SO2 + 2C

First, look for the amount of substance of sulfur:
n(S) = m / M
n(S) = 14.8 g/32 g / mol = 0.4625 mol
n(CO) = m (CO) / M (CO)
M(CO) = 12 + 16 = 28 g/mol
n(CO) = 19.9 g/28 g/mol = 0.71 mol
S in excess, so for calculating we take CO:
n(SO2) = n(CO)/2 = 0.71 mol/2 = 0.355 mol
m(SO2) = M(SO2)*n(SO2)
M(SO2) = 32 + 16*2 = 64 g/mol
m(SO2) = 64 g/mol * 0.355 mol = 22.74 g